Night Plan UT 02-Dec-2020

Sunset: 17:19

Twilight Ends 18:40
  • Starting Configuration is Bino-PEPSI
  • PEPSI Pol Checkout
    • AGW3/AGW4 checkout and alignment beginning at sunset (1 hr) - critical
    • Observations of Polarimetric Standard calibrator gamma equ (21:10:20 +10:07:53 - hits 30deg limit at 21hr local time) - 0.5 hr
  • Reconfigure for Binocular LUCI - authorize LUCI-1
    • OSCO (NCPA off and NCPA on)
    • Authorize Bino-LUCI
      • ESM/L1 and SL/L2 observations of UDS z~10 target - Rothberg (3hrs) - clear or thin cirrus - DIMM 1.2" or less
        • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/UDS_z10_ESM_Science
        • Observations are mixed-mode: L1-ESM and L2-SL. Same setup on both sides (N3.75, G210 HKspec, CWL=1.6, 0.75" slit)
        • Blind offset along slit - center on AO Ref first then offset to faint galaxy - target is H=25.3 (won't see in acq images). AO Ref star remains in slit during observations
        • 51 degrees from Moon
    • If PEPSI and LBC do not need additional time
      • ESM observations of Palomar 2 - (2hrs with standards, 1hr without standards) - may punt this target if PEPSI and/or LBC need time
        • if photometric observe P247U before and after target (cycles through all 5 broad band filters)
        • 10 degrees from the Moon
      • ESM observations of M67
        • highest priority is R=16.3 AO Ref star - verify if same issues occurred as first E/D Night
  • If PEPSI needs more time
    • 3 hours for further AGW3 and AGW4 work on IT 7545 which ended softly (medium-high)
    • 2 hours for AGW3 and AGW4 alignment (medium)
  • If no PEPSI time needed, and LBC needs time
    • Focus offsets (2hrs)
    • On-sky Gain/Linearity measurements (using commissioning methods) (2hrs). This requires photometric skies.
    • Continue TMS testing if needed (or passive TMS data collection)
  • Poor Weather/Poor Seeing Condition Programs
    • ESM data collection M34 (3 filters) , M52 (3 filters and bright and faint AO Ref star), M67 (3 filters and bright and faint AO Ref star) - if you can close the loop take the data
    • ICRF3 Spectroscopy (Rothberg)
      • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/Poor_Weather/ICRF3_Rothberg_2020B
      • Observe telluric first
      • Double Check high Dec objects to make sure tellurc/science are close spatially
    • If clear but poor seeing, please observe standard stars, cycling through z,J,H,K,Ks
      • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/Poor_Weather/LUCI_Phot_ZP
      • Select stars to use based on visibility - OT file is in directory, export the scripts as needed
      • Please observe the same star at least two times at different airmasses

Night Summary

PEPSIPOL installation went long (insufficient staff, staff doing tasks they weren't familiar with, failed glycol connector that had to be replaced, power breaker on C-ring for SX POL was off, etc.). And winds rose above limits about the end of evening twilight.

04:30 UT Status: Both AGWs are connected and operating, Both Polarimeters are connected and communicating, still have science fibers remaining to connect to polarimeters.

05:30 UT Status: Installation of science fibers complete. Ready for on-sky if weather will cooperate.

Details

Configured and authorized for PEPSIPOL / PEPSIPOL

05:52 UT - Preparing to Open

05:59 UT - Open with wind gusting to 18 m/s, sky mostly clear

06:05 UT - Stop and Restart GCSs to get images in ACQUIRE preset

06:16 UT - Both Active Optics are converged - as much as you can converge in 2.2 - 3 arcsec bubbly seeing

06:20 UT - Target stars (RSP2011 137) are in the fiber and pinhole guiding is underway - as much as you can in 2.2 - 3 arcsec bubbly seeing

06:30 UT - Trefoil saturated at 5000 nm on DX. Clear Active Optics on DX M1. Besides horrible seeing, the pinhole guiding is shoving the wfs spots into the left side of the lenslet apertures. Per IT 7703, the GCS guide star position limit is set to 2.54 arcsec radius rather than the usual 0.54 arcsec radius. This prevent wandering in the pinhole guiding from killing the off-axis guide loop.

06:45 UT - New target ( ) repeated preset 3-times.

06:49 UT - New target at EL=43 deg. These are real targets, but we are mostly testing acquisition and active optics. Seeing was momentarily better down here at 43 deg.

06:57 UT - New target to start a science integration. (AO guys aren't willing to fight us for the telescope in 2.5 arcsec seeing.)

07:04 UT - Cleared Active Optics on DX M1 again, image was getting elliptical.

07:40 UT - New Target

-- BarryRothberg - 01 Dec 2020
Topic revision: r3 - 02 Dec 2020, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback